Hernan works in the same restaurant and has been working for a longer time.
4. Herman works in the same restaurant and has been working for longer time. In 1 hour, he can make 6 pizzas or 14 burgers. (5 points) Table 1: productivity, quantities of pizza and bur r hour JudieHemon Pizza (per hour) 4 Burger (per hour) 14 a. How many burgers are equal to 1 pizza? b. How many pizzas are equal to 1 burger? c. How much is Hermans opportunity cost of pizza [Note that 1 hour can be used to make either d. e. pizzas or burgers, not both]? How much is Hermans opportunity cost of burger? Who has comparative advantage in burger, pizza? What determines comparative advantage?
Answer #1

a ) Ans:

For Herman , 1 pizza = 14 / 6 = 2.33 burgers

For Judie , 1 pizza = 8 / 4 = 2 burgers

b ) Ans:

For Herman , 1 burger = 6 / 14 = 0.42 pizza

For Judie, 1 burger = 4 / 8 = 0.5 pizza

c ) Ans:

Herman's opportunity cost of pizza = 14 / 6 = 2.33

d ) Ans:

Herman's opportunity cost of burger = 6 /14 = 0.42

e ) Ans:

Herman has comparative advantage in burger and Judie has comparative advantage in pizza.

Lower opportunity cost of producing a good determines comparative advantage.

Explanation:

Herman has comparative advantage in burger due to lower opportunity cost for producing it as compared to Judie. Whereas Judie has comparative advantage in pizza due to lower opportunity cost for producing it as compared to Herman.